# Simple Yet Effective: Extracting Private Data Across Clients in Federated Fine-Tuning of Large Language Models

### v1

@@ 4.2.2 Frequency-Prioritized Prefix Sampling (FP Sampling)
which defines subsets of prefixes that occur exactly $\sigma$ times. This induces a frequency-based decomposition of the unique prefixes in $\text{SUP}(P_{c})$:
@@
Given a frequency threshold $\sigma_{a}$, we define the set of frequent prefixes as:
$P_{f\geq\sigma_{a}}=\bigcup_{\sigma\geq\sigma_{a}}P_{\sigma},$ (5)
which is then sorted in descending order of frequency. Notably, setting $\sigma_{a}=1$ yields the full set of generalized contextual prefixes, $P_{f\geq 1}=\text{Set}(\text{SUP}(P_{c}))$, sorted in descending order of frequency.
Given a prefix budget $B$, we construct the final prefix set by selecting the top-$B$ prefixes from $P_{f\geq\sigma_{a}}$, thereby prioritizing high-frequency prefixes during sampling.
4.2.3 Latent Association Fine-tuning (LAFt)
We hypothesize that a model’s susceptibility to extraction attacks stems from its latent ability to associate two underlying conceptual distributions: (1) $\mathcal{A}$ — the distribution over prefixes that are semantically or syntactically likely to precede PII; and (2) $\mathcal{B}$ — the distribution over actual PII instances. Let $\text{Dist}(\mathcal{A},\mathcal{B};\theta)$ denote the degree of association between these two distributions under model parameters $\theta$.
Because this association is implicitly encoded in the model’s internal representations, we propose a method to reduce this distance through parameter updates—a technique we term Latent Association Fine-tuning (LAFt). The core idea is to fine-tune the model to minimize $\text{Dist}(\mathcal{A},\mathcal{B};\theta)$, thereby reinforcing its internal linkage between indicative prefixes and corresponding PII, ultimately improving its capacity for PII extraction.
To implement this, we construct a fine-tuning dataset $D_{\text{ft}}$ by pairing frequent prefixes with known PII instances from the attacker’s dataset:
@@ 4.2.3 Latent Association Fine-tuning (LAFt)
where $P_{f}$ is the set of frequent prefixes derived from $D_{a}$, and $S_{a}$ is the attacker’s known PII set. We then fine-tune the model using the standard causal language modeling objective:
@@
The updated model $\theta^{\prime}$ is then used for extraction, using prefixes from either $P_{f}$ or $P_{c}$.
@@ 5.1 Experimental Setup
We detail the experimental setup and implementation to ensure reproducibility.
@@
We simulate a federated system with 5 clients using a label-skewed non-IID data partitioning method based on clustering of language embeddings (Li et al., 2023). Each client receives a comparable number of samples. Federated fine-tuning is conducted on legal tasks using the OpenFedLLM framework Ye et al. (2024b), with FedAVG as the aggregation method over 10 communication rounds. All clients adopt parameter-efficient fine-tuning (LoRA) and a shared prompt template. Hyperparameter settings and implementation details are detailed in Appendix H.
After federated fine-tuning, we evaluate the utility of the final global model on a held-out global test set. Following common practice, we compare this with a centrally trained (non-FL) baseline on the same test set. Results are shown in Table 8.
@@
We use QWen1-8B (Bai et al., 2023) and Baichuan2-7B (Yang et al., 2023), both pre-trained primarily on Chinese corpora. Model performance is evaluated with Coverage Rate (CR), Efficiency (EF), and Victim-exclusive Extracted PII (VxPII, defined as $|(S_{v}\setminus S_{a})\cap S_{E}|$), defined to measure extraction accuracy and completeness.
@@
We designate client 0 as the attacker and client 1 as the victim, and evaluate three strategies: (1) PII-contextual prefix sampling. The attacker builds a prefix set $P_{c}$ from its local dataset $D_{0}$ with prefix length $\lambda=50$. Each prefix queries the global model 15 times, generating up to $m=10$ tokens per query—sufficient to recover most PIIs with manageable cost. (2) Frequency-prioritized sampling. Prefixes in $\text{Set}(\text{SUP}(P_{c}))$ are ranked by frequency to form $P_{f\geq 1}$ and used in descending order. Sweeping the prefix budget $B$ (base-10 exponential scale) varies the frequency threshold $\sigma_{a}$, enabling analysis of coverage–efficiency trade-offs. (3) Latent association fine-tuning. The attacker fine-tunes the global model (1 epoch, LR = 5e-5, LoRA: $r=16$, $\alpha=32$) using 10k frequent prefixes and 10k randomly sampled PIIs from its own data to reinforce prefix–PII associations.
Further implementation details are provided in Appendix H.3.2.
@@
To ensure fair evaluation, we define the set of victim-exclusive PIIs as $(S_{v}\setminus S_{a})$ and apply two filters: (1) only PIIs $s_{i}\in S_{v}$ that do not appear in the attacker’s training corpus ($s_{i}\notin U_{a}$) are retained; and (2) we enforce a longest common prefix constraint to eliminate interference between PIIs—e.g., those sharing prefixes—which may confuse the determination of which PII has been extracted (see Equation (1)):
@@
Metrics are computed on this filtered and prefix-disjoint set, as defined in Equations (2)–(3).
5.2 Main Results
5.2.1 RQ1: To what extent can such attacks extract PII at maximum?
We begin by evaluating the maximum coverage rate (CR) achievable by our proposed attack methods. Specifically, we use the full set of PII-contextual prefixes $P_{c}$ to query the federated fine-tuned LLMs. To explore the upper bound of extraction capability, we also employ the sorted generalized contextual prefix set, $\text{Set}(\text{SUP}(P_{c}))=P_{f\geq 1}$, though this is limited to the Qwen1-8B model due to its high computational cost.
Table 1 reports the results. Using the PII-contextual prefix set $P_{c}$, we achieve a CR of 22.93% on Qwen1-8B and 28.95% on Baichuan2-7B. Applying Latent Association Fine-tuning (LAFt) improves Qwen1-8B’s CR to 28.30%, though the corresponding efficiency (EF) remains low at 0.24%, indicating a high proportion of irrelevant or duplicated outputs. The count of unique, victim-exclusive PII (VxPII) remains relatively small.
Notably, leveraging the broader prefix set $P_{f\geq 1}$ substantially increases Qwen1-8B’s CR to 56.57%. However, this comes at the expense of extremely low efficiency (0.01%), suggesting that while coverage increases, most samples are not useful.
For Baichuan2-7B, applying LAFt does not lead to a direct improvement in CR, but we later show that LAFt consistently recovers a distinct set of PII instances.
Table 1: Attack performance of client 0 (attacker) against client 1 (victim) using different methods. The $(S_{v}\setminus S_{a})$ contains 8870 unique victim-exclusive PII items.
Model Prefix Set LAFt CR EF VxPII Count Prefix Set Size
Qwen1-8B $P_{c}$ without 22.93% 0.1910% 2034 71006
$P_{c}$ with 28.30% 0.2357% 2510 71006
$P_{f\geq 1}$ without 56.20% 0.0110% 4985 3013161
$P_{f\geq 1}$ with 56.57% 0.0111% 5018 3013161
Baichuan2-7B $P_{c}$ without 28.95% 0.2411% 2568 71006
$P_{c}$ with 28.46% 0.2370% 2524 71006
5.2.2 RQ2: What is the distribution of extracted PII?
Figure 2 presents the label distribution of deduplicated victim-exclusive PII instances extracted by the Qwen1-8B and Baichuan2-7B models. The horizontal bar plot displays label names on the y-axis and the corresponding instance counts on the x-axis.
Figure 2: Label distribution of deduplicated victim-exclusive PII instances.
@@ 5.2.2 RQ2: What is the distribution of extracted PII?
Among all extracted PII types, "Address", "Birthday", and "Name" appear most frequently. In contrast, categories such as "Work Experience" and "Work Place" occur less often but remain notable. More complex categories, such as "Medication Record", are not extracted at all. This absence is primarily due to the nature of the extraction attack and the evaluation protocol, which only considers model-generated PII that exactly matches the ground truth. Since complex PII types often manifest as long free-text segments, it is difficult for models to reproduce them with verbatim accuracy.
5.2.3 RQ3: How Effective Are FP Sampling and LAFt in Extracting VxPII?
We present a Venn diagram of the VxPII (Victim-exclusive extracted PII) sets produced by different methods in Figure 3. As shown, the Frequency-Prioritized Prefix Sampling (FP Sampling) and the Latent Association Fine-tuning (LAFt) methods identify overlapping but distinct sets of PIIs, , suggesting that each method captures different aspects of information.
We also analyze the VxPII distribution produced by the LAFted model and find that, although LAFt uncovers a different subset of instances, the overall label distribution remains similar across models. This indicates that LAFt tends to uncover new instances of already-well extracted PII types, rather than broadening the diversity of PII types. This phenomenon may be attributed to the way we construct the LAFt fine-tuning dataset $D_{\text{ft}}$, where frequent prefixes are paired with random PIIs. To encourage the discovery of rarer PII types, future work might explore pairing prefixes with less frequent or underrepresented PII types during fine-tuning.
Figure 3: Venn diagrams showing the overlap of VxPII sets extracted by different methods. (a) Comparison of PII-contextual prefix set $P_{c}$ vs. frequency-prioritized prefix set $P_{f\geq 1}$ with a prefix budget of $B=10000$, on Qwen1-8B without LAFt; (b) Comparison of VxPII extracted by Qwen1-8B with and without LAFt, using the full $P_{c}$ prefix set; (c) Comparison of VxPII extracted by Baichuan2-7B with and without LAFt, using the full $P_{c}$ prefix set.
5.2.4 RQ4: What Is the Tradeoff Between Coverage and Efficiency?
Figure 4 illustrates the tradeoff between coverage rate (CR) and efficiency (EF) as the prefix budget $B$ increases. Specifically, CR rises significantly while EF declines sharply for both prefix sets $P_{c}$ and $P_{f\geq 1}$. As more PII are extracted, discovering new, non-duplicate PII becomes increasingly difficult. This observation points to a potential tradeoff decision in real-world attack scenarios, where expanding the budget could lead to diminishing efficiency gains.
@@ 5.2.4 RQ4: What Is the Tradeoff Between Coverage and Efficiency?
Figure 4: Coverage rate (CR) and efficiency (EF) under varying prefix budgets $B$ for prefix sets $P_{c}$ and $P_{f\geq 1}$. The budget values are scaled exponentially (base 10).
5.3 Ablation Study
To evaluate the robustness of our PII extraction approach, we conduct a cross-validation experiment by iteratively designating each client as the attacker while treating the remaining clients as victims. This helps ensure that our method is not biased towards a particular client. As shown in Table 9, the proposed extraction attack demonstrates consistent effectiveness across all attacker–victim pairings.
5.4 Evaluating the Effectiveness of Simple PII Masking
@@
We evaluate the effectiveness of a simple data sanitization strategy that masks PII using existing annotations. Each PII instance in the training data is replaced with a string of asterisks (*) of equal length. We then re-fine-tune FedLLM on this sanitized dataset and perform the PII-contextual Prefix Sampling attack once again. Table 2 compares the attack performance with and without the PII masking defense.
@@
As shown in the results, the number of extracted VxPII is only slightly reduced. To understand this, we examine the document frequency of VxPII instances—that is, how often each appears in the training data. Figure 7 shows that PII masking significantly lowers the frequency of most VxPII, suggesting that our annotations effectively cover the majority of PII. Interestingly, some VxPII with zero document frequency—indicating they were absent from the masked dataset—were still extracted.
Based on these observations, we identify two main reasons for the limited effectiveness of the masking defense: (1) Pretraining data contamination: Our training data, sourced from publicly available legal documents, likely overlaps with the pretraining corpora of models like Qwen1-8B and Baichuan2-7B. (2) Incomplete PII labeling: Some PII instances may remain unlabeled—and therefore unmasked—in the training data. In the realistic scenarios, attackers can customize PII definitions, making exhaustive and comprehensive coverage inherently challenging.
Pretraining data contamination is difficult to eliminate, as LLM providers rarely disclose their pretraining corpora. To mitigate this, we adopt a conservative strategy: we compare the VxPII extracted from FedLLM ($F$) with that from its un-fine-tuned base model ($B$), and subtract $B$ to isolate PII memorized during federated fine-tuning. Table 3 shows that even after removing $B$, a substantial number of VxPII remain in $F\setminus B$, confirming memorization during fine-tuning. Figure 8 further shows that $F\setminus B$ exhibits a similar label distribution of VxPII as observed in Figure 2, supporting the validity of our method.

### v2

@@ 4.2.2 Frequency-Prioritized Prefix Sampling
This yields
@@
Given a threshold $\sigma_{a}$, the frequent prefix set is
$P_{f\geq\sigma_{a}}=\bigcup_{\sigma\geq\sigma_{a}}P_{\sigma},$
sorted by frequency. Setting $\sigma_{a}=1$ recovers the full contextual prefix set. With a budget $B$, we select the top-$B$ prefixes from $P_{f\geq\sigma_{a}}$, thereby emphasizing frequent contexts.
4.2.3 Latent Association Fine-tuning
We conjecture that a model’s vulnerability to PII extraction stems from its capacity to capture the conditional probability $\text{P}(\mathcal{B}\mid\mathcal{A};\theta)$ under model parameters $\theta$, where $\mathcal{A}$ denotes prefixes that typically precede PIIs and $\mathcal{B}$ represents the corresponding PII instances.
Since the association between PII and their prefixes is implicitly encoded in the model’s representations, we propose Latent Association Fine-tuning (LAFt), which updates parameters to maximize $\text{P}(\mathcal{B}\mid\mathcal{A};\theta)$. The goal is to strengthen the mapping between indicative prefixes and PII, thereby improving extraction.
As the first step, we build a fine-tuning dataset $D_{\text{ft}}$ by pairing frequent prefixes with known PII:
@@ 4.2.3 Latent Association Fine-tuning
where $P_{f}$ is the frequent-prefix set from $D_{a}$, and $S_{a}$ the attacker’s PII set. The model is then fine-tuned with the standard causal LM objective:
@@
The updated model $\theta^{\prime}$ is subsequently used for extraction with prefixes from $P_{f}$ or $P_{c}$.
To note that, LAFt follows the semi-honest FL setting such that the fine-tuned model remains local and is not uploaded to the server.
@@ 5.1 Experimental Setup
Our federated setup consists of two main components: the data partitioning across clients and the federated fine-tuning procedure.
For the data partitioning, we simulate a system with 5 clients using a label-skewed non-IID partitioning based on clustering of language embeddings (Li et al., 2023), and ensured that each client receives a comparable number of samples.
For the federated fine-tuning, we perform training on legal tasks using the OpenFedLLM framework (Ye et al., 2024b), with FedAVG (McMahan et al., 2023) as the aggregation method over 10 communication rounds. All clients adopt parameter-efficient fine-tuning (LoRA) and a shared prompt template. Hyperparameter settings and implementation details are provided in Appendix K.
After the federated fine-tuning, we evaluate the utility of the final global model on a held-out global test set. Following common practice, we compare it to a centrally trained (non-FL) baseline evaluated on the same test set. The results are reported in Table 8 in the Appendix.
@@
Since our data and tasks are derived from Chinese legal documents, we focus primarily on LLMs with proficiency in Chinese. Specifically, we evaluate Qwen1-8B (Bai et al., 2023), Baichuan2-7B (Yang et al., 2023), Qwen3-8B (Yang et al., 2025), GLM4-8B (GLM et al., 2024), and Llama3-Chinese (Rayrtfr, 2024). [^3] All models are publicly available on HuggingFace: Qwen1, Baichuan2, Qwen3, GLM4, Llama3-Chinese.
We evaluate model performance using two primary metrics: Coverage Rate (CR), Efficiency (EF). In addition, we introduce an intermediate metric, Victim-exclusive Extracted PII (VxPII), defined as $|(S_{v}\setminus S_{a})\cap S_{E}|$, which directly measures the amount of extracted information.
@@
We designate client 0 as the attacker and client 1 as the victim, and evaluate three strategies: (1) PII-contextual prefix sampling. The attacker builds a prefix set $P_{c}$ from its local dataset $D_{0}$ with prefix length $\lambda=50$. Each prefix queries the global model 15 times, generating up to $m=10$ tokens per query—sufficient to recover most PIIs with manageable cost. (2) Frequency-prioritized sampling. Prefixes in $\text{Set}(\text{SUP}(P_{c}))$ are ranked by frequency to form $P_{f\geq 1}$ and used in descending order. Sweeping the prefix budget $B$ varies the frequency threshold $\sigma_{a}$, enabling analysis of coverage–efficiency trade-offs. (3) Latent association fine-tuning. The attacker fine-tunes the global model (1 epoch, LR = 5e-5, LoRA: $r=16$, $\alpha=32$) using 10k frequent prefixes and 10k randomly sampled PIIs from its own data to reinforce prefix–PII associations. Further implementation details are provided in Appendix K.3.2.
@@
To ensure a fair evaluation, the set of victim-exclusive PIIs $(S_{v}\setminus S_{a})$ is obtained by applying two filters: (1) retain only those victim PIIs that do not appear in the attacker’s training corpus (i.e., $s_{i}\in S_{v}$ but $s_{i}\notin U_{a}$); and (2) remove PIIs that share a common prefix to avoid ambiguity in identifying which PII was extracted (see Equation (1)). This is enforced by constraining the length of the longest common prefix (LCP) between any two PIIs:
@@
Metrics defined in Equations (2) and (3) are then computed on this filtered, prefix-disjoint set.
5.2 Results and Discussions
RQ1: How effective is the PII extraction attack using contextual prefixes?
We first evaluate the coverage rate (CR) and efficiency (EF) of our extraction attacks by querying federated fine-tuned LLMs using the PII-contextual prefix set $P_{c}$. Table 1 presents the results. With $P_{c}$, our attack achieves a considerable CR of 22.93% on Qwen1-8B and 28.95% on Baichuan2-7B.
Table 1: Summary of attack results using the PII-contextual prefix sampling method (with and without LAFt), where client 0 (attacker) targets client 1 (victim). The victim-exclusive set $(S_{v}\setminus S_{a})$ includes 8,870 unique PII items.
Model Prefix Set CR EF VxPII Count Prefix Set Size
wo LAFt
Qwen1-8B $P_{c}$ 22.93% 0.1910% 2034 71006
Baichuan2-7B $P_{c}$ 28.95% 0.2411% 2568 71006
Qwen3-8B $P_{c}$ 30.69% 0.2556% 2722 71006
GLM4-9B $P_{c}$ 28.20% 0.2348% 2501 71006
Llama3-Chinese-8B $P_{c}$ 19.73% 0.1643% 1750 71006
Qwen1-8B $\mathrm{Set}(\mathrm{SUP}(P_{c}))$ 56.20% 0.0110% 4985 3013161
Baichuan2-7B $\mathrm{Set}(\mathrm{SUP}(P_{c}))$ 53.56% 0.0105% 4751 3013161
w LAFt
Qwen1-8B $P_{c}$ 28.30% 0.2357% 2510 71006
Baichuan2-7B $P_{c}$ 28.46% 0.2370% 2524 71006
Qwen1-8B $\mathrm{Set}(\mathrm{SUP}(P_{c}))$ 56.57% 0.0111% 5018 3013161
Baichuan2-7B $\mathrm{Set}(\mathrm{SUP}(P_{c}))$ 52.16% 0.0102% 4627 3013161
Figure 3: Label distribution of deduplicated victim-exclusive PII extracted by Qwen1-8B (without LAFt, using prefix set $P_{c}$). Results for Baichuan2-7B are shown in Appendix Figure 11.
To understand what types of PII are most vulnerable, we analyze the extracted instances. Figure 3 shows the label distribution of deduplicated victim-exclusive PII extracted by Qwen1-8B (without LAFt). The results for Baichuan2-7B are provided in Appendix Figure 11.
@@ 5.2 Results and Discussions
The most frequently extracted PII categories include "Address", "Birthday", and "Name", while others such as "Work Experience" and "Work Place" occur less often but remain notable. More complex types like "Medication Record" are not extracted at all. This is primarily due to the evaluation protocol, which only credits model outputs that match ground truth exactly. Complex PII often appears as long free-text spans, making verbatim reproduction difficult.
@@
Figure 4: Coverage rate (CR) and efficiency (EF) under varying prefix budgets $B$ for prefix sets $P_{c}$ and $P_{f\geq 1}$. Prefix set $P_{f\geq 1}$ is frequency-sorted in descending order (see Section 4.2.2). Budget values are scaled exponentially (base 10); model used is Qwen1-8B.
To estimate an upper bound of extraction capability, we evaluate with the generalized prefix set $\mathrm{Set}(\mathrm{SUP}(P_{c}))$, which includes all potential contextual prefixes. As shown in Table 1, expanding $P_{c}$ to $\mathrm{Set}(\mathrm{SUP}(P_{c}))$ increases CR to 56.57% (Qwen1-8B) and 53.56% (Baichuan2-7B). However, this gain comes at a steep cost in efficiency—dropping EF to only 0.01%—indicating most queries yield redundant or irrelevant content.
We further investigate this CR–EF tradeoff in Figure 4, which illustrates how CR and EF vary with prefix budget $B$ for prefix sets $P_{c}$ and $P_{f\geq 1}$. As $B$ increases, CR improves, but EF declines sharply. This suggests diminishing returns in efficiency when scaling up the number of queries to discover new PII instances.
RQ2: How effective is frequency-prioritized prefix sampling?
As shown in Figure 5, frequency-prioritized (FP) sampling does not extract more VxPII instances than the contextual prefix set $P_{c}$, contrary to our hypothesis in Section 4.2.2. This result suggests that the contextual cues embedded in $P_{c}$ are already strong indicators of LLM memorization, and that memorization cannot be inferred solely from co-occurrence frequency. Instead, it likely arises from more complex interactions between corpus semantics, model architecture, and pre-training dynamics.
Figure 5: VxPII counts under varying prefix budgets ($B$) for prefix sets $P_{c}$ and $P_{f\geq 1}$. Prefix set $P_{f\geq 1}$ is frequency-sorted in descending order (see Section 4.2.2) and truncated to match the size of $P_{c}$ here.
Despite this, FP sampling captures highly distinct subsets of memorized PII. As shown in Figure 6(a), the Venn diagram comparison reveals that 49.9% of the VxPII extracted by FP sampling on Qwen1-8B and 65.02% on Baichuan2-7B are not discovered by the $P_{c}$ method. This highlights FP sampling’s complementary strength in uncovering diverse memorized content.
RQ3: How effective is Latent Association Fine-tuning?
As shown in Table 1, applying Latent Association Fine-tuning (LAFt) significantly improves the CR of Qwen1-8B by 5.37%, raising it to 28.30%, and increases EF to 0.24%, indicating enhanced extraction performance. For Baichuan2-7B, LAFt does not yield a direct improvement in CR, but, as depicted in Figure 6(b), it facilitates the identification of additional distinct PII instances.
These results demonstrate that LAFt is an effective method for increasing the diversity of extracted PII, complementing the FP sampling approach. The extent of the improvement achieved by LAFt is influenced by the construction of the fine-tuning dataset $D_{\text{ft}}$ and the choice of hyperparameters. In this study, we adopt a consistent setting by constructing $D_{\text{ft}}$ through pairing frequent prefixes with randomly sampled PII and fine-tuning the model for one epoch to ensure a fair comparison. However, further exploration of personalized strategies—tailored to models with different architectures and pre-training conditions—could potentially yield better performance.
Figure 6: Venn diagrams showing overlap between VxPII sets extracted by different methods. (a) Comparison of VxPII sets using PII-contextual prefixes $P_{c}$ vs. frequency-prioritized prefixes $P_{f\geq 1}$ at prefix budget B=10, [^0] B=10{,}000 (without LAFt). (b) Comparison of VxPII sets extracted with and without LAFt on Qwen1-8B and Baichuan2-7B using the full $P_{c}$ prefix set.
5.3 Cross-Client Evaluation of Extraction Robustness
To assess the robustness of our PII extraction method across different clients, we perform a cross-client evaluation where each client is iteratively designated as the attacker, while the remaining clients act as victims. This setup ensures that the extraction performance is not biased toward any particular client.
As shown in Table 2, our method achieves consistently high coverage rates across all attacker–victim pairings, demonstrating its generalizability and effectiveness in diverse settings.
Table 2: Coverage rates (CR) of extraction attacks across different attacker–victim client pairs with a prefix budget $B=10000$. Prefixes are randomly sampled from each attacker’s corresponding set $P_{c}$. “–” indicates self-attack scenarios, which are not applicable.
Attacker ID Victim ID
0 1 2 3 4
0 - 10.91% 12.89% 10.93% 11.88%
1 11.97% - 12.41% 11.46% 12.35%
2 12.56% 11.39% - 11.65% 12.74%
3 12.07% 10.82% 12.04% - 11.99%
4 12.26% 11.36% 13.25% 11.21% -
5.4 PII Sanitization Defense
@@
We evaluate the effectiveness of a simple data sanitization strategy that masks PII using existing annotations. Each PII instance in the training data is replaced with an equal-length sequence of asterisks (*). We then re-fine-tune FedLLM on this sanitized dataset and reapply the PII-contextual Prefix Sampling attack. Table 3 compares the attack performance with and without the PII masking defense.
@@
The results show only a slight reduction in the number of extracted VxPII. To investigate this, we analyze the document frequency of extracted VxPII—that is, how often each appears in the training corpus. Figure 8 demonstrates that masking substantially lowers the frequency of most VxPII, suggesting that our annotations successfully cover the majority of PII. Interestingly, some VxPII with zero document frequency—absent from the sanitized dataset—were still extracted.
Based on these observations, we attribute the limited effectiveness of masking to two factors. First, pretraining data contamination: our training data, drawn from publicly available legal documents, likely overlaps with the pretraining corpora of models such as Qwen1-8B and Baichuan2-7B. Second, incomplete PII labeling: some PII instances may be missing from annotations, and in practice attackers can redefine PII categories, making exhaustive coverage fundamentally difficult.
5.5 Disentangling the Effect of Data Contamination
Pretraining data contamination is difficult to avoid, as LLM providers rarely disclose their pretraining corpora. To mitigate this influence, we adopt two strategies:
Using open-source LLMs.
To eliminate interference from pretraining memorization, we use open-source or semi-open-source LLMs that are not pretrained on Chinese legal documents: OLMo2-7B (OLMo et al., 2024) and Llama2-7B (Touvron et al., 2023). As shown in Table 4, after just one round of federated fine-tuning, the CR exceeded 23%, achieving performance comparable to or better than Qwen1-8B and Baichuan2-7B. This confirms that our findings are not simply an artifact of pretraining contamination.
Table 4: Extraction results on non-Chinese LLMs after a single round of federated fine-tuning.
Model Prefix Set CR EF
OLMo2-1124-7B $P_{c}$ 23.04% 0.1919%
Llama-2-7B $P_{c}$ 26.41% 0.2200%
Subtracting contaminated memorization.
For Chinese-proficient LLMs that may contain contamination, we adopt a subtraction-based approach. Specifically, we compare the VxPII extracted from the fine-tuned FedLLM ($F$) with those from its base model ($B$), and compute $F\setminus B$ to isolate PII memorized during federated fine-tuning. Table 5 shows that even after subtracting $B$, a substantial number of VxPII remain in $F\setminus B$, confirming memorization during fine-tuning. Furthermore, Figure 10 demonstrates that $F\setminus B$ exhibits a distribution of VxPII labels similar to Figure 3, supporting the validity of this analysis.